Metallica Bearings, Inc., is a young start-up company. No dividends will be paid on the stock over the next 8 years because the firm needs to plow back its earnings to fuel growth. The company will pay a dividend of $12 per share 9 years from today and will increase the dividend by 6 percent per year thereafter.

  

If the required return on this stock is 13 percent, what is the current share price?
Answer #1

Solution:-

Price 8 years from now = D9 / (Required return - Growth rate)

                                             =12 / (0.13 - 0.06)

                                             = 12 / 0.07

                                             = 171.43

Current share price =D8*Present value of discounting factor(rate%,time period)

                                          =$171.43 / 1.13^8

                                         = $171.43 / 2.658

                                         =$64.50(Approx.)